If f(z) is an entire function such that f(z)=f(-z) then there exists an entire function g(z) such that f(z)=g(z^2)


1
Expert's answer
2021-11-22T14:48:01-0500

If a complex function is analytic at all finite points of the complex plane , then it is said to be entire.


let

"z=x+iy"

"f(z)=u(x,y)+iv(x,y)"

since "f(z)=f(-z)" :

"u(x,y)=u(-x,-y),v(x,y)=v(-x,-y)"


then:

since "z^2=(x+iy)^2=(x^2-y^2)+ixy" , then:


"g(z^2)=u_1(x^2-y^2,xy)+iv_1(x^2-y^2,xy)"


since "f(z)=g(z^2)" , then:


"u_1(x^2-y^2,xy)=u(x,y)=u(-x,-y)"


"v_1(x^2-y^2,xy)=v(x,y)=v(-x,-y)"